OpenStudy (anonymous):

Find two consecutive integers whose sum is 22.

ganeshie8 (ganeshie8):

trial and error

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

consecutive integers u want ? 9 + 10 = 19 REJECT, its less than 22

ganeshie8 (ganeshie8):

10+11 = 21 REJECT, still less than 22

ganeshie8 (ganeshie8):

11+12=23 REJECT, more than 22

OpenStudy (anonymous):

Find two consecutive odd integers whose difference is 2.

ganeshie8 (ganeshie8):

so did we miss 22 ?

OpenStudy (anonymous):

i copy it wrong

OpenStudy (anonymous):

lol

ganeshie8 (ganeshie8):

ohhk :) do trial and error for this also pick any two consecutive odd integers and test

OpenStudy (anonymous):

Infinite solutions, the difference between two odd numbers is always 2

ganeshie8 (ganeshie8):

Yup ! good job !!